OpenStudy (anonymous):

y=(tan(x))^arcsin(x) , whats dy/x ? Wont this be d(tan(x)^arcsin(x))/dtan(x) * d(tan(x))/dx according to the chain rule ? Where am I going wrong?

OpenStudy (anonymous):

to simply this take ln on both sides

OpenStudy (anonymous):

\[y=\tan(x)^{\arcsin(x)}\]

OpenStudy (anonymous):

ln y = arcsinx (ln tanx)

OpenStudy (anonymous):

What's wrong with what I said about the chain rule ?

OpenStudy (anonymous):

ur doing it wrong..its complicated..if u want it that way u treat one fn as constant and then diff and add it to the derivative keeping second fn constant

OpenStudy (anonymous):

like if uv got f(x) and g(x) and y=f^g dy/dx = gf^(g-1) + (lnf)f^g

OpenStudy (anonymous):

Getting confused. Dont understand why I need to add functions here.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

now u get it?

OpenStudy (anonymous):

dy/dx = gf^(g-1) + (lnf)f^g makes it to add ..

OpenStudy (anonymous):

but can you tell me whats wrong with the chain rule ? am afraid I have not understood the chain rule itself properly cause of it got wrong.

OpenStudy (anonymous):

f(x)=h(g(x)) , f'(x) = h'(g(x))*g'(x) , is this not what I did too ?

OpenStudy (anonymous):

derivative of outside function w.r.t its argument * derivative of argument w.r.t independent variable

OpenStudy (anonymous):

u know it right..but here we cannot use the chain rule

OpenStudy (anonymous):

u tell me...r u in class 12?

OpenStudy (anonymous):

yes am in class 12

OpenStudy (anonymous):

why cant I use chain rule here ?

OpenStudy (anonymous):

for using the chain rule for sin(e^x) do u need to know the derivative of sinx or not?

OpenStudy (anonymous):

yes I know

OpenStudy (anonymous):

ur thinking of simpler derivatives for the chain rule..for eg we know a^x ka derivative...so hencewe can derive a^f(x)

OpenStudy (anonymous):

wont that be cos(e^x) * e^x ?

OpenStudy (anonymous):

yes...just an example..the thing is that there is no standard derivative of tanx ^x

OpenStudy (anonymous):

ur treating ur function as f(g(x)) where f is tanx and g is arcsinx which is wrong

OpenStudy (anonymous):

yes am treating it like that , why is it wrong ??

OpenStudy (anonymous):

agar aisa hota to it wldve been tan (arcsinx) or something

OpenStudy (anonymous):

acha , see this , how do you differenciate x^x w.r.t x then ?

OpenStudy (anonymous):

cz remember in a composite fn in place of X there is another function, here theres no function in place of x itself na? do u c?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so u do logarithmic differentiation or u maro style and keep one fn constant and do that pellet...right...?

OpenStudy (anonymous):

thing is I have not yet come to logarithmic differentiation abhi tak ... all I know is chain rule

OpenStudy (anonymous):

I can blindly convince myself that I might need logarithmic differentiation for this , but I truly dont understand why chain rule cant do.

OpenStudy (anonymous):

this is only an extension of chain rule..bt ive taught u now havent i whenever u c exponents like y=f^g take log on both sides and then u can apply chain rule

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is a chain rule question...just gotta use brains and log properties

OpenStudy (anonymous):

is it like when g and f are 2 functions of x and h is another function = g^f , you need to use logarithmic ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

see y = x^sinx ln y= sinxlnx now use chain rule

OpenStudy (anonymous):

r u in delhi??

OpenStudy (anonymous):

thanks a lot..u gave me something to think abt...mazaa aagaya..bt dont think m nerdy...

OpenStudy (anonymous):

yes .. ;) lol no problems , what do you do ?

OpenStudy (anonymous):

ive given iit right one month back...

OpenStudy (anonymous):

hey him1618, i think chain rule can suffice ... no log needed

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Like you were in 12th and gave IIT or just gave IIT being elder?

OpenStudy (anonymous):

abhi 12th pass ki hai yyar..beizzati mat karo

OpenStudy (anonymous):

i passed in 2011..gave iit in 2011

OpenStudy (anonymous):

lol you are brainy

OpenStudy (anonymous):

fine fine wrote AIEEE?

OpenStudy (anonymous):

thanx...aakash sahib kahan gaye...

OpenStudy (anonymous):

yeah..but marks nahi lagaye..wanted to be free

OpenStudy (anonymous):

Aakash went to temple .. will be right back soon he says.

OpenStudy (anonymous):

Are you on any IM ? What's your gtalk ?

OpenStudy (anonymous):

facebook...u in delhi?

OpenStudy (anonymous):

yes I am in Delhi.

OpenStudy (anonymous):

You ?

OpenStudy (anonymous):

wt school?

OpenStudy (anonymous):

Springdales

OpenStudy (anonymous):

you ?

OpenStudy (anonymous):

k.....add Himanish...studied at St. Columba's School

OpenStudy (anonymous):

ur pretty...g2g...ill add u

OpenStudy (anonymous):

thanks haha :) , sure.

OpenStudy (anonymous):

hey him... sorry .. now i realize my mistake .. u r rite ... we need to find the derivative of x^arcsin(x) and then multiply with diff of tan (x^arcsin(x))

OpenStudy (anonymous):

we find the first one using log .. as u said

OpenStudy (anonymous):

dont pste ur fb id on public pages....aakash....sk goyal ke phatte aise hi nhi maare..bye

OpenStudy (anonymous):

hun?

OpenStudy (anonymous):

Damn posted my fb profile here !

OpenStudy (anonymous):

toldya..b careful next tym

OpenStudy (anonymous):

r u a quizzer?

OpenStudy (anonymous):

quizzer ?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

mean do u go fr quizzes and all?

OpenStudy (anonymous):

No , you ?

OpenStudy (anonymous):

yeah..i am..just inquiring whether uve heard of the columban open quiz...nvm...ive sent u d request

OpenStudy (anonymous):

what request ?

OpenStudy (anonymous):

btw what are you planning to study next after 12th ?

OpenStudy (anonymous):

engineering physics btech

OpenStudy (anonymous):

Oh , which university ?

OpenStudy (anonymous):

abhi result to aane do yaar..waise i think ill get into one of the IITs

OpenStudy (anonymous):

very nice.

OpenStudy (anonymous):

Was IIT easy this year ?

OpenStudy (anonymous):

How many hours do you study for IIT ? Went to any coaching ?

OpenStudy (anonymous):

yes..only i screwed up...

OpenStudy (anonymous):

narayana...no defined hours of study...

OpenStudy (anonymous):

went to any coaching classes at narayana or you stay at narayana ?

OpenStudy (anonymous):

assuming you mean "narayna" the area ..

OpenStudy (anonymous):

narayana iit academy

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

hav u added?

OpenStudy (anonymous):

Oh , wait i'll add now.

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

how do you get an area under a function f(x) between spans x=a and x=b when you integrate it ?

OpenStudy (anonymous):

How come it fetches you the area ? intuition behind it ?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

No ?

OpenStudy (anonymous):

u divide the area under the graph into small strips along the x-axis, each having width dx

OpenStudy (anonymous):

Dude , gimme an IM. I cant login to facebook right now , my sis's logged in and if I log out , she'll scream , its her laptop. Give me your google talk or msn or yahoo or AIM or skype or whatever.

OpenStudy (anonymous):

these rectangular strips have height f(x) at point x, and width dx no IM now...my bits is on monday..sm1 sees m dead...